OpenStudy (anonymous):

solve 5(2x+1)=3(5x-4)

jhonyy9 (jhonyy9):

10x+5=15x-12 17=10x x=17/10

OpenStudy (saifoo.khan):

Crrct.

OpenStudy (anonymous):

saifoo instead of saying correct, give a medal

OpenStudy (saifoo.khan):

lol, a Medal for a Medal.

OpenStudy (anonymous):

10x+5 = 15x - 12 17 = 5x x = 17/5

OpenStudy (anonymous):

5(2x+1)=3(5x-4) expanding both sides... 10x + 5 = 15x -12 adding 12 to both sides and subtracting 10x from both sides... 17 = 5x dividing both by 5 ... 17/5 = x